An individual was employed as a church custodian for a fixed salary of $30,000 per year. This salary was payment for a normal 40 hour work week. The custodian was responsible for cleaning and maintaining the church and church property. The church also had a grade school that they operated. The grade school was separated from the church, but the church employed all the teachers and staff in the school. The school was looking for a part-time janitor. The school wanted to pay the janitor $9.00 per hour for five (5) hours of work each school day evening.

The church custodian applied for, and was given the job as the part-time janitor for the school. He was to continue his day job as the custodian for the church.

My questions are as follows: How should the custodian's pay be calculated? Can the two jobs be separated for payroll purposes? What types of issues are involved given your understanding of the Fair Labor Standards Act as it is described in Chapter 2?
Answer #1

1. The custodian's pay will be for an year

- From church - $30,000p.a

- From School - $16,425p.a

$9 * 5 hours * 365 = $16,425p.a

Per hour rate

- From church - $14.45

- From School - $9

Calculation for church pay per hour

$30000/52 weeks =577 per week

$577/40 hours a week = 14.45per hour

2. The church and School even though seperated but managed by same Management Committee. Thus it can be considered same for calculations of payroll purposes.

3. Provisions regarding Fair Labour Standard Act :

All employees that hold positions determined to be covered under the mandatory overtime provisions of the FLSA are covered. Overtime-eligible employees must be compensated with overtime pay or compensatory time for all hours worked over 40 in a single workweek. For all hours worked in excess of 40 during each work week, employees will receive overtime at the rate of one and one-half times the employee’s regular rate

Application in the question :

The church pays $14.45 per hour whereas for overtime work in school it pays $9 which is lesser than one time of normal pay. Thus we can say that Church Management is violating provisions of Fair Labour Standard Act